holly rioux invested some money at 3.5% simple interest, and $5000 more than 3 times this amount at 4%. In one year, she earned $1440 in interest. How much did she invest at each rate?
Answer by rfer(16322)

.035x+.04(3x+5000)=1440
.035x+.12x+200=1440
.155x=1240
x=$8000
3x+5000=$29000